OpenStudy (anonymous):

Question 1.1. Simplify. If x = –4, evaluate –x. (Points : 1) 8 4 0 –4 Question 2.2. Simplify. 6 – [–(–1)] (Points : 1) –7 7 –5 5 Question 3.3. Simplify. | –3.9| + | 3.9| (Points : 1) –7.8 0 7.8 8 Question 4.4. Evaluate the expression if d = –3 |d| (Points : 1) –3 0 1 3 Question 5.5. Evaluate the expression. |18 – 4| (Points : 1) 22 14 –22 –14

OpenStudy (anonymous):

METAL and FAN

OpenStudy (anonymous):

1) Is x= -4 then -x=-(-4), -x=? 2) 6-[-(-1)] = 6-[+1]= 6-1=? 3) |-3.9| + |3.9| = +3.9 + +3.9 = 3.9+3.9 = ? 4) |d|= |-3|, |d|=+3 5)|18-4| = |14| = +14 |a| or the absolute value of 'a' is just the positive value of the variable 'a'.

OpenStudy (anonymous):

i dont get what your saying @amlim14

OpenStudy (anonymous):

Hahahaha. I'm saying that when you see those two vertical bars enclosing any number or letter, it means the absolute value, which also means just the positive (+) value of that number or letter.

OpenStudy (anonymous):

so can u help me with the answers @amlim14

OpenStudy (anonymous):

The answers are already here with solutions. 1.1) Is x= -4 then -x=-(-4), -x=? 2.2) 6-[-(-1)] 6-[+1] 6-1=? 3.3) |-3.9| + |3.9| = +3.9 + +3.9 = 3.9+3.9 = ? 4.4) |d|= |-3|, |d|=+3 5.5)|18-4| = |14| = +14

OpenStudy (anonymous):

ugh i dont get it... can u like do this 1 is 4 2 is 5 lol like that cause im stupid lol

OpenStudy (anonymous):

@amlim14

OpenStudy (anonymous):

anyone hellpppp please!!!

OpenStudy (anonymous):

NVM I GET IT

OpenStudy (anonymous):

THANK YOU VERYYY MUCH